A manufacturing firm just received a shipment of 15 assembly parts, of slightly varied sizes, from a vendor. The manager knows that there are only 12 parts in the shipment that would be suitable. He examines these parts one at a time.

a. Find the probability that the first part is suitable. (Round your answer to 2 decimal places.)

Probability:_______


b. If the first part is suitable, find the probability that the second part is also suitable. (Round your answer to 4 decimal places.)

Probability:_______


c. If the first part is suitable, find the probability that the second part is not suitable. (Round your answer to 4 decimal places.)

Probability:_______
Answer #1

Total number of parts = 15

Number of suitable parts = 12

Number of non suitable parts = 15 - 12 = 3

a) P(first part is suitable) = Number of suitable parts / Total number of parts

= 12/15

= 0.80

b) If first part is suitable, number of parts remaining is 14, of which 11 are suitable.

P(second part is suitable | first part is suitable) = 11/14

= 0.7857

c) If first part is suitable, number of parts remaining is 14, of which 3 are not suitable.

P(second part is not suitable | first part is suitable) = 3/14

= 0.2143